# Quantum Abduction: A New Paradigm for Reasoning under Uncertainty

### v1

@@ Abstract
Abductive reasoning—the search for plausible explanations—has long been central to human inquiry, from forensics to medicine and scientific discovery. Yet formal approaches in AI have largely reduced abduction to eliminative search: hypotheses are treated as mutually exclusive, evaluated against consistency constraints or probability updates, and pruned until a single “best” explanation remains. This reductionist framing overlooks the way human reasoners sustain multiple explanatory lines in suspension, navigate contradictions, and generate novel syntheses. This paper introduces quantum abduction, a non-classical paradigm that models hypotheses in superposition, allows them to interfere constructively or destructively, and collapses only when coherence with evidence is reached. Grounded in quantum cognition and implemented with modern NLP embeddings and generative AI, the framework supports dynamic synthesis rather than premature elimination. Case studies span historical mysteries (Ludwig II of Bavaria, the “Monster of Florence”), literary demonstrations (Murder on the Orient Express), medical diagnosis, and scientific theory change. Across these domains, quantum abduction proves more faithful to the constructive and multifaceted nature of human reasoning, while offering a pathway toward expressive and transparent AI reasoning systems.
@@ 1 Introduction
Motivation
Many critical domains—criminal investigation, clinical diagnostics, scientific discovery—require reasoning under ambiguity, contradiction, and evolving evidence. Classical abduction, as formalized in logic, Bayesian inference, or set-covering models, typically reduces to eliminative selection: hypotheses compete within a fixed space until one prevails. This does not reflect the way human experts actually reason, often maintaining competing explanations in play until evidence forces convergence.
Contributions
We propose quantum abduction, a new framework where hypotheses are held in superposition, interact through interference, and collapse contextually into dominant or hybrid syntheses. The contributions of this paper are fourfold: (i) a conceptual model grounded in both quantum cognition and computational semantics; (ii) implementation guidelines bridging symbolic clarity with statistical power; (iii) cross-domain demonstrations, from forensic reasoning to scientific theory change; and (iv) a comparative assessment against classical abductive frameworks. The present paper focuses on conceptualization, comparative analysis, and design guidelines. Full formal and computational development—encompassing Hilbert-space formulations, interference operators, and integration with subsymbolic methods—is reserved for subsequent work.
Structure of the Paper
Section 2 reviews abductive reasoning traditions and prior work in quantum cognition and entangled heuristics. Section 3 outlines the computational premises of quantum abduction, motivates their practical value, and sketches implementation pathways. Sections 4.14.4 present case studies across forensic, literary, medical, and scientific domains. Section 5 situates our model within related work. We conclude in Section 6 with implications and future directions.
@@ 2 Background
Classical Abduction
Abductive reasoning—variously termed abduction, abductive inference, or retroduction—is a form of logical inference that proposes the most plausible explanation for a given set of observations. The concept was first articulated by the American philosopher and logician Charles Sanders Peirce [13] in the late 19th century, who framed abduction alongside deduction and induction as one of the three fundamental modes of reasoning. More recent accounts emphasize its cognitive grounding and role as a general reasoning pattern, both in scientific discovery and in everyday inference [9, 20].
Unlike deduction, which guarantees the truth of its conclusions if the premises are true, abduction yields hypotheses that are tentative and provisional. Its conclusions are best understood as plausible guesses or informed conjectures, often described in terms such as “best available” or “most likely” explanations. Abduction also differs from induction: whereas induction generalizes from particular cases to broader regularities, abduction remains tied to specific observations, advancing hypotheses without claiming universality.
From the 1980s onward, the notion of abduction attracted growing interest in applied fields including law, clinical practice, computer science, and artificial intelligence. Diagnostic expert systems, for example, often used abduction to generate candidate explanations for observed symptoms or evidence. However, these implementations revealed both practical and theoretical limits. Practically, they were constrained by the computational resources of the time and by the lack of robust learning mechanisms. Theoretically, they were limited by the eliminative character of classical abduction itself: hypotheses are typically treated as mutually exclusive, and the reasoning process progresses by discarding inconsistent candidates until a single “best” explanation remains.
This eliminativist dynamic, grounded in the principles of non-contradiction and excluded middle, creates brittleness precisely in the kinds of situations where human experts exhibit flexibility—namely, in contexts of contradictory, incomplete, or entangled evidence. Classical abduction struggles to sustain multiple explanatory threads in suspension, forcing early closure at the expense of richer, hybrid accounts. These shortcomings provide the rationale for exploring quantum abduction, a framework that models inference in terms of superposition and interference, allowing explanatory alternatives to coexist and interact before eventual resolution.
Quantum Cognition
@@
This shift enables the modeling of cognitive processes in which contradictions, ambiguities, and contextual dependencies are not merely tolerated but are constitutive of reasoning itself. Rather than forcing premature resolution, quantum cognition allows multiple possibilities to coexist until a contextual “measurement” induces collapse.
For our purposes, the relevance of quantum cognition lies in its demonstration that quantum-inspired formalisms can provide tractable, predictive accounts of human reasoning where classical models fail. While the ontological claims of quantum mechanics do not apply—no one is suggesting that abductive inference literally “creates” reality—the formal machinery of Hilbert spaces, superposition, and interference offers a powerful metaphor and computational toolkit. These tools allow us to construct a novel paradigm of quantum abduction, in which hypotheses are sustained as overlapping amplitudes, shaped by mutual reinforcement or contradiction, and resolved only when constraints induce collapse into a dominant or hybrid explanation.
Entangled Heuristics and Strategic Inference
The work entitled “Entangled Heuristics for Agent-Augmented Strategic Reasoning” ([5]) introduced a hybrid deductive reasoning architecture in which strategic heuristics are activated and composed through semantic entanglement. In that context, heuristics extracted from strategic texts (e.g., military treatises, corporate frameworks) are treated not as discrete rules but as semantically interdependent potentials, with transformer models employed to measure and exploit these semantic interdependencies [4]. Conflicting or complementary heuristics are not eliminated but fused, inducing the emergence of novel and conceptually richer heuristic syntheses. This process embodies a quasi-ontological creativity: new strategic concepts arise through heuristic entanglement rather than being selected from a fixed menu of pre-existing options.
While the philosophical stance of that project was deliberately constructivist—treating inference as creation—quantum abduction maintains a stronger epistemic orientation. Here, hypotheses represent real but uncertain explanatory elements about a determinate world. The entanglement mechanism remains valuable, however: embedding hypotheses in a Hilbert-style semantic space, modeling their interferences, and enabling synthetic recombination provides the computational machinery needed to capture inference dynamics in uncertain contexts.
@@ 3 Computational Premises for Quantum Abduction
Classical abduction, as first articulated by Peirce, has often been formalized in artificial intelligence as eliminative search or best-explanation selection within a fixed hypothesis set. This formalization rests on the law of non-contradiction and the principle of the excluded middle: given contradictory hypotheses, one must be rejected, and reasoning proceeds until a single “winning” explanation remains. The brittleness of this approach becomes evident precisely in the kinds of complex contexts where human experts are able to keep multiple explanatory lines “in suspense,” allowing contradictory cues to coexist until further constraints emerge.
Quantum abduction refrains from claiming that reasoning itself creates reality, as in the ontological collapse of quantum mechanics, nor does it pursue the constructive synthesis of new strategic concepts, as in [5]. In this respect, it shares with classical abduction the tenet that inference unfolds within a determinate world, where in fact only one sequence of events has occurred. Yet it departs radically from the classical tradition by sustaining hypotheses in superposition. This superposition is epistemic rather than ontological: it will ultimately “collapse” when an agent’s explanatory state, constrained by accumulating evidence, converges toward a (possibly composite) account that best aligns with reality. The strength of the quantum metaphor lies not in importing physical indeterminacy, but in providing a computational formalism that avoids the premature eliminativism of classical logic.
With this caveat, the relationship with quantum mechanics remains, formally and computationally, substantial; even more so with respect to synthesis from conceptual entanglement as in [5]. The mathematical machinery of quantum theory—Hilbert spaces, projection operators, and interference dynamics—offers a powerful representational framework for abductive reasoning. Superposition here means that alternative explanations can coexist as weighted amplitudes in a shared semantic space; interference models how evidence can simultaneously reinforce some explanatory combinations while attenuating others; and collapse denotes the emergence of a coherent explanatory synthesis once evidential constraints cross a threshold of adequacy. In this way, as represented in Figure 1, quantum abduction differentiates itself from classical approaches by preserving the richness and heterogeneity of the hypothesis space while still ensuring eventual convergence on a single determinate reality.
Minimal formal sketch
@@
$\Delta\alpha_{i}\;\propto\;\max_{j}\,\langle H_{i}\mid O_{j}\rangle,\qquad\langle H_{i}\mid O_{j}\rangle\approx\frac{\mathbf{h}_{i}\cdot\mathbf{o}_{j}}{\|\mathbf{h}_{i}\|\,\|\mathbf{o}_{j}\|}.$
Inter-hypothesis influence is captured by an interference matrix $I_{ij}$ (constructive/destructive), so updates combine evidence and interactions:
$\alpha_{i}^{(t+1)}\;\propto\;\alpha_{i}^{(t)}+\eta\Big(\underbrace{\sum_{j}\langle H_{i}\mid O_{j}\rangle}_{\text{evidence}}\;+\;\underbrace{\sum_{k\neq i}I_{ik}\,\alpha_{k}^{(t)}}_{\text{interference}}\Big).$
Collapse occurs when a coherence functional exceeds a threshold (or when a decision must be made); collapse may select a single $H_{i}$ or a hybrid synthesis.
Classical Abduction (Elimination)Observations Hypothesis Set $\{H_{1},H_{2},H_{3}\}$ Eliminate Inconsistent Hypotheses Final Hypothesis $H^{*}$Quantum Abduction (Superposition & Interference)Observations Superposed Hypotheses $\alpha_{1}|H_{1}\rangle+\alpha_{2}|H_{2}\rangle+\alpha_{3}|H_{3}\rangle$ Interference & Amplitude Update Epistemic Collapse to Coherent Explanation
Figure 1: Classical abduction prunes alternatives through eliminative search. Quantum abduction keeps hypotheses in epistemic superposition, allows interference, and collapses only when evidence suffices.

### v2

@@ Abstract
Abductive reasoning—the search for plausible explanations—has long been central to human inquiry, from forensics to medicine and scientific discovery. Yet formal approaches in AI have largely reduced abduction to eliminative search: hypotheses are treated as mutually exclusive, evaluated against consistency constraints or probability updates, and pruned until a single “best” explanation remains. This reductionist framing fails on two critical fronts. First, it overlooks how human reasoners naturally sustain multiple explanatory lines in suspension, navigate contradictions, and generate novel syntheses. Second, when applied to complex investigations in legal or scientific domains, it forces destructive competition between hypotheses that later prove compatible or even synergistic, as demonstrated by historical cases in physics, astronomy, and geology. This paper introduces quantum abduction, a non-classical paradigm that models hypotheses in superposition, allowing them to interfere constructively or destructively, and collapses only when coherence with evidence is reached. Grounded in quantum cognition and implemented with modern NLP embeddings and generative AI, the framework supports dynamic synthesis rather than premature elimination. For immediate decisions, it models expert cognitive processes; for extended investigations, it transforms competition into “co-opetition” where competing hypotheses strengthen each other. Case studies span historical mysteries (Ludwig II of Bavaria, the “Monster of Florence”), literary demonstrations (Murder on the Orient Express), medical diagnosis, and scientific theory change. Across these domains, quantum abduction proves more faithful to the constructive and multifaceted nature of human reasoning, while offering a pathway toward expressive and transparent AI reasoning systems.
arXiv Note: This version corresponds to the final published article in Sci 2025, Vol. 7, Issue 4, Article 182. DOI: 10.3390/sci1010000. Formatted for arXiv compliance by removing publisher-specific formatting while preserving all scholarly content, corrections, figures, and references.
@@ 1 Introduction
1.1 Motivation
Many critical domains—criminal investigation, clinical diagnostics, scientific discovery—require reasoning under ambiguity, contradiction, and evolving evidence. Classical abduction, as formalized in logic, Bayesian inference, or set-covering models, is typically reduced to eliminative selection: hypotheses compete within a fixed space until one prevails. This reductionist approach fails on two distinct timescales.
First, in immediate decision contexts (emergency medicine, real-time diagnostics), it does not reflect how human experts actually reason, often maintaining competing explanations in play until evidence forces convergence. Second, in extended investigative contexts (complex criminal cases, scientific theory development), the eliminative paradigm creates counterproductive competition among research teams, investigative units, or theoretical schools, when collaborative “co-opetition”—where competing hypotheses inform and strengthen each other’s development—would be more fruitful.
1.2 Dual Role of Quantum Abduction
Quantum abduction addresses both challenges through a unified framework. For immediate decisions, it models the parallel processing observed in expert cognition, where multiple explanatory threads remain active until collapse is necessary. For extended investigations, it transforms competitive hypothesis testing into a coordinated exploration where:
  • Different investigative teams pursue distinct but informationally entangled hypotheses
  • Evidence gathered for one hypothesis can constructively or destructively interfere with others
  • The framework acts as a meta-cognitive coordinator, tracking how partial insights from competing approaches might synthesize into richer explanations
  • Resources are allocated based on current amplitude distributions rather than premature winner-takes-all decisions
This “co-opetition” model is particularly valuable in complex criminal investigations (where tunnel vision on a single suspect can blind investigators to alternative leads) and scientific research (where premature paradigm commitment can delay breakthrough syntheses for decades).
1.3 Nature and Originality of This Work
This paper presents a research framework with initial conceptual validation for a novel approach to abductive reasoning. The framework, termed quantum abduction, is entirely original in its specific formulation, though it builds upon established work in quantum cognition [1] and extends our prior research on entangled heuristics [2]. While the latter focuses on strategic concept creation through ontological synthesis, quantum abduction is epistemically grounded—it seeks the best explanation of a single determinate reality rather than creating new conceptual ontologies. The use of quantum formalism as a computational metaphor for cognitive processes, while not new, is applied here in an original way to the specific problem of abductive inference.
From a computational and formal standpoint, we adopt the Hilbert space formalism from quantum cognition to model interference, contextuality, and order effects in explanation; see Section 3.1 for the full rationale and limits.
1.4 Structure of the Paper
The remainder of the paper is organized as follows. Section 2 motivates the use of Hilbert space formalisms for abductive inference and clarifies the scope and limits of the quantum analogy. Section 3 introduces the quantum abduction framework and its computational components. Section 4 illustrates its use across forensic, clinical, literary, and scientific contexts. Section 5 provides an initial empirical assessment using a minimal benchmark. Section 6 contrasts the approach with logic-based, Bayesian, set-covering, and other alternative abductive traditions. Section 7 concludes with implications, limitations, and directions for future development within the broader research program. A glossary of key terms is provided in Appendix A, and formal specifications are consolidated in Appendix B.
@@ 2 Background
2.1 Classical Abduction
Abductive reasoning—variously termed abduction, abductive inference, or retroduc-tion—is a form of logical inference that proposes the most plausible explanation for a given set of observations. The concept was first articulated by the American philosopher and logician Charles Sanders Peirce [3] in the late 19th century, who framed abduction alongside deduction and induction as one of the three fundamental modes of reasoning. More recent accounts emphasize its cognitive grounding and role as a general reasoning pattern, both in scientific discovery and in everyday inference [4, 5].
Unlike deduction, which guarantees the truth of its conclusions if the premises are true, abduction yields hypotheses that are tentative and provisional. Its conclusions are best understood as plausible guesses or informed conjectures, often described as “best available” or “most likely” explanations. Abduction also differs from induction: whereas induction generalizes from particular cases to broader regularities, abduction remains tied to specific observations, advancing hypotheses without claiming universality.
From the 1980s onward, the concept of abduction has garnered increasing interest in various applied fields, including law, clinical practice, computer science, and artificial intelligence. Diagnostic expert systems, for example, often used abduction to generate candidate explanations for observed symptoms or evidence. However, these implementations revealed both practical and theoretical limits. Practically, they were constrained by the computational resources of the time and by the lack of robust learning mechanisms. Theoretically, they were limited by the eliminative character of classical abduction itself: hypotheses are typically treated as mutually exclusive, and the reasoning process progresses by discarding inconsistent candidates until a single “best” explanation remains.
This eliminativist dynamic, grounded in the principles of non-contradiction and the excluded middle, creates brittleness precisely in situations where human experts exhibit flexibility—namely, in contexts of contradictory, incomplete, or entangled evidence. Classical abduction struggles to sustain multiple explanatory threads in suspension, forcing early closure at the expense of richer, hybrid accounts. These shortcomings provide the rationale for exploring quantum abduction, a framework that models inference in terms of superposition and interference, allowing explanatory alternatives to coexist and interact before eventual resolution.
2.2 Quantum Cognition
@@
We use these tools as a modeling formalism for non-classical aspects of reasoning (interference, contextuality), not as a claim about quantum hardware or ontological indeterminacy; detailed clarifications appear in Section 3.1.
2.3 Entangled Heuristics and Strategic Inference
The work entitled “Entangled Heuristics for Agent-Augmented Strategic Reasoning” [2] introduces a hybrid deductive reasoning architecture in which strategic heuristics are activated and composed through semantic entanglement. In that context, heuristics extracted from strategic texts (e.g., military treatises, corporate frameworks) are treated not as discrete rules but as semantically interdependent potentials, with transformer models employed to measure and exploit these semantic interdependencies [7]. Conflicting or complementary heuristics are not eliminated but rather fused, leading to the emergence of novel and conceptually richer heuristic syntheses. This process embodies a quasi-ontological creativity: new strategic concepts arise through heuristic entanglement rather than being selected from a fixed menu of pre-existing options.
While the philosophical stance of that project was deliberately constructivist—treating inference as a form of creation—quantum abduction maintains a stronger epistemic orientation. Here, hypotheses represent real but uncertain explanatory elements about a determinate world. The entanglement mechanism remains valuable; however, embedding hypotheses in a Hilbert-style semantic space, modeling their interferences, and enabling synthetic recombination provide the computational machinery needed to capture inference dynamics in uncertain contexts.
@@ 2.3 Entangled Heuristics and Strategic Inference
Figure LABEL:fig:pipeline_comparison anticipates Section 3 by contrasting eliminative abduction with superposition interference.
@@ 3 Computational Premises for Quantum Abduction
Classical abduction, as outlined in Section 2, typically assumes a fixed hypothesis set and an eliminative search process: contradictions are resolved by discarding candidates until a single explanation remains. Rather than reiterating these limitations, we now turn to the computational consequences of moving beyond them.
Quantum abduction shares the classical commitment to a determinate underlying reality—only one sequence of events has actually occurred—but departs from eliminativism in its representation of uncertainty. Hypotheses are maintained in an epistemic superposition, with amplitudes that evolve as evidence accumulates. Observations act as projections in a semantic Hilbert space; interactions among hypotheses are modeled through an interference matrix that captures reinforcement, contradiction, or semantic overlap. Collapse marks explanatory convergence, which may result in a dominant or hybrid explanation rather than a single survivor from a fixed menu.
This section introduces the operational machinery that enables these dynamics, centered on the mathematical summary in Box 1. We treat Hilbert space tools—vectors, projections, interference—not as ontological commitments but as a computational formalism that resists premature elimination and preserves information until evidence warrants decisive convergence. Full derivations appear in Appendix B.
3.1 Why Quantum?—Interpretational Clarifications
As anticipated in the Introduction, “quantum” is used here in an epistemic and computational sense rather than an ontological one. Section 2 gave the conceptual motivation; here we state the implications at a more formal level and show how they guide the computational design.
Three Distinct Notions
  • Quantum cognition: Hilbert space structure to model belief states and their context-sensitive evolution (this paper).
    • Intuition: A Hilbert space offers a geometric setting in which each hypothesis is a vector and each observation acts as a projection. Inner products encode explanatory alignment; interference terms capture how alternatives can reinforce or attenuate one another. This yields contextuality and order effects providing the non-classical compositionality needed for abductive reasoning under contradiction.
  • Quantum computation: algorithms on quantum hardware (not used here).
  • Classical simulation of quantum-like dynamics: classical hardware implementing interference/superposition mathematics (our regime).
Why the Formalism Matters Here
The formalism captures three features of explanatory reasoning that classical abduction struggles with:
  1. 1.
    Interference of explanatory alternatives, yielding non-classical updates (violations of the law of total probability when hypotheses interact);
  2. 2.
    Contextuality, where the plausibility of a hypothesis depends on which other hypotheses are co-considered;
  3. 3.
    Order effects in evidence evaluation, as documented in cognitive experiments [6].
Superposition and collapse are epistemic: coexistence of competing explanations prior to convergence; collapse as emergence of a coherent synthesis under sufficient constraints.
Relation to Entangled Heuristics
Methodologically, we share machinery with Entangled Heuristics for Agent-Augmented Strategic Reasoning [2] (transformers encode semantic interdependence; interference drives composition). The interpretational aim differs:
  • In [2], semantic superposition supports ontological innovation (new strategic constructs).
  • Here, superposition tracks epistemic uncertainty about a determinate reality.
Thus, the same tools (semantic embeddings, transformer-mediated similarity, interference dynamics) serve ontological synthesis there, and epistemic inference here.
@@ 3.1 Why Quantum?—Interpretational Clarifications
Transformers and LLMs (Conceptual Primer)
Hypotheses and observations are expressed in natural language; we therefore need a representation that preserves semantic structure. Transformer encoders map text to high-dimensional vectors, where semantic proximity corresponds to geometric proximity. This does not entail that transformers “understand” the hypotheses; rather, they supply a coordinate system in which superposition and interference can be computed. LLMs, in turn, articulate hybrid explanations in natural language; they do not decide scores.
3.2 Transformer and LLM Integration as Epistemic Infrastructure
Embedding Model
We use the transformer model Sentence-BERT [8] (e.g., all-MiniLM-L6-v2, d=384d{=}384) to encode hypotheses $H_{i}$ and observations $O_{j}$ into a shared semantic space: $\mathbf{h}_{i}=\mathrm{SBERT}(\text{text}(H_{i}))$ and $\mathbf{o}_{j}=\mathrm{SBERT}(\text{text}(O_{j}))$. Results are largely invariant to the specific SBERT variant up to rotation; higher-dimensional models increase fidelity at modest cost.
Clarifying the Objects Involved
Each hypothesis $H_{i}$ and observation $O_{j}$ begins as a text string. These are not numerical objects. To use them computationally, we convert them into fixed-length numerical vectors using the SBERT encoder.
What SBERT Does (Intuitive Explanation)
SBERT is a transformer-based model that maps any piece of text—ranging from a single word to a short paragraph—into a vector in $\mathbb{R}^{384}$. All embeddings have the same dimensionality regardless of the input length. This uniformity is why cosine similarity between $\mathbf{h}_{i}$ and $\mathbf{o}_{j}$ is well-defined.
In principle, long hypotheses could be decomposed into clause- or sentence-level embeddings and then pooled (e.g., by weighted averaging or attention). For this initial framework we adopt the simpler and widely used “one SBERT vector per text” strategy, leaving fine-grained composition to future work.
Meaning of $H_{i}$, $\mathbf{h}_{i}$, and $e_{i}$
To avoid ambiguity:
  • $H_{i}$ = hypothesis as a linguistic description;
  • $\mathbf{h}_{i}$ = its embedding in $\mathbb{R}^{384}$;
  • $O_{j}$ = observation as text;
  • $\mathbf{o}_{j}$ = its embedding in $\mathbb{R}^{384}$.
The projection score for hypothesis $i$ at time $t$ is
$e_{i}^{(t)}=\cos(\mathbf{h}_{i},\mathbf{o}^{(t)}).$
It has a single index because it measures the activation of one hypothesis, even though its computation involves two embeddings.
Qualitative vs. Quantitative Projection Matrices
The tables in Section 4 use qualitative marks (v = supports, x = contradicts) purely for readability. In the actual computation, each entry is replaced by a cosine similarity between SBERT embeddings, optionally modulated by an interference coefficient $\kappa_{ij}$ described below.
Projection (Evidence Activation)
Similarity $e_{i}^{(t)}=\cos(\mathbf{h}_{i},\mathbf{o}_{j})$ acts as a projection operator, activating or suppressing amplitudes without forcing single-winner selection.
Interference Estimation
We set $I_{ij}=\kappa_{ij}\cdot\text{sim}(H_{i},H_{j})$ with cosine similarity over SBERT vectors. $\kappa_{ij}\in[-1,1]$ encodes exclusivity/complementarity and can be initialized (i) heuristically from domain priors, (ii) via weak supervision (optimize ECS/HQI on solved cases), or (iii) by expert elicitation smoothed through metric learning.
Amplitude Update (Single Step)
Input: alpha[1..n], I[1..n,1..n], evidence e[1..n], eta in (0,1]
for i in 1..n:
[…]
    * alpha[k] )
Normalize: alpha = alpha_tilde / ||alpha_tilde||_2
Collapse if max_i |alpha[i]|ˆ2 > tau; else continue
Mix Operator and LLM Articulation
When multiple hypotheses retain high amplitudes and exhibit constructive interference, the system prepares a candidate hybrid by combining their embeddings via a weighted semantic blend. The precise operator, its justification, and its role in case studies are elaborated in Section 3.3; at this point we note only that the LLM’s role is strictly limited to articulating the hybrid in natural language, not to scoring or selecting it.
Human-in-the-Loop
Experts adjust $\kappa_{ij}$ (conflict/compatibility), calibrate collapse thresholds $\tau$, and edit LLM phrasing for clarity and faithfulness. This design follows a human-in-the-loop principle [9], in which domain experts operate at a semantic rather than mathematical level. They do not need Hilbert space or amplitude calculus; their inputs are judgments about complementarity, conflict, or partial overlap among hypotheses (i.e., signs and strengths for $\kappa_{ij}$, and acceptable collapse thresholds $\tau$). Within a broader centaurian paradigm [10], the assistant provides formal control and interpretability over superposition/interference dynamics while preserving human agency over meaning and decision.
Full definitions, equations, and derivations are provided in Appendix B; here we retain only the operational summary.
3.3 The Mix Operator for Hypothesis Synthesis
When several hypotheses retain high amplitudes and exhibit constructive interference, the system applies the mix operator to generate hybrid explanations:
$\text{mix}(H_{i},H_{j})=\lambda_{ij}\mathbf{h}_{i}+(1-\lambda_{ij})\mathbf{h}_{j},\qquad\lambda_{ij}=\frac{|\alpha_{i}|^{2}}{|\alpha_{i}|^{2}+|\alpha_{j}|^{2}}.$ (1)
The natural-language articulation follows the LLM strategy above. Formal variants and derivations appear in Appendix B.